# Train Once, Deploy Anywhere: Realize Data-Efficient Dynamic Object Manipulation

### v1

@@ V-D2 Analysis on Instance Segmentation
0.05 0.92 0.96 0.79 0.74
0.10 0.90 0.97 0.80 0.71
0.25 0.81 0.92 0.73 0.48
0.50 0.00 0.00 0.00 0.00
V-D3 Analysis on Visual Servo tracking
In GEM, we utilize the visual servo tracking module to convert the dynamic object manipulation problem into static object manipulation in the view of the recurrent probabilistic network. This study analyzes the tracking stability and velocity upper limit in real-world manipulation. Specifically, we adjust the conveyor belt to various speeds and make the robot hand track a moving object based on the visual servo tracking module. It is observed that the maximum stable tracking speed in simulation is approximately 0.25 m/s, and it reaches about 0.3 m/s in the real world. In Fig. 17, we visualize the dynamic curves of the control target signal and corresponding end-effector position. We can find that the end-effector approaches the moving object fast and tracks it stably under the velocity of 0.2 m/s. However, the end-effector fails to track the object when the velocity is 0.6 m/s due to its speed limit.
Afterwards, we validate the real-world success rates of GEM in the settings of different moving speeds. The results are reported in Table IV. Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results, we can observe that the success rates are similar when the conveyor belt runs at a low speed. When the speed is 0.25 m/s, the success rates drop due to stricter manipulation precision requirements. When the speed arrives at 0.50 m/s, the robot hand cannot perform any manipulation task because it cannot keep pace with the movement of the target object.
TABLE V: Analysis on the effect of the memory encoder.
@@ V-D3 Analysis on Visual Servo tracking
$\times$ 0.72 0.60 0.69 0.55
$\checkmark$ 0.87 0.78 0.85 0.74
V-D4 Analysis on Memory Modeling
As discussed in Section III-D, enabling a network to access historical memory is important for reducing action prediction ambiguity. To this end, we devise the memory encoder shown in Fig. 8. In experiments, we find that without this encoder, the robot arm could oscillate at the initial position of manipulation without approaching the target, leading to task failure. This phenomenon is more frequent when a single network learns multiple skills. This is because the policy fails to distinguish whether the manipulation operation has not started or has been completed. Our designed memory encoder can effectively address this problem by enabling the network to know which actions have been performed. We compare the performances of GEM with and without this memory encoder and report the results in Table V. According to the results, we can observe that the memory encoder significantly boosts success rates.
To more thoroughly analyze how long a sequence the memory encoder can memorize, we design a toy experiment of reciting a random digit sequence. Specifically, we update the structure of this memory encoder to support accepting a single digit as input and predicting the next digit per step. Then, we randomly generate a sequence of $n$ digits and train the memory encoder to recite this sequence. After training, we test how many digits can be correctly recited. After repeating this experiment by setting different values of $n$, we find that the memory encoder can correctly recite a very long sequence like 1000 digits after sufficient epochs of training. Such a memory capability is strong enough for existing embodied policies, because current manipulation tasks usually involve at most tens of times of policy inference.
@@ V-D4 Analysis on Memory Modeling
Direct 0.81 0.73 0.79 0.67
Probabilistic 0.87 0.78 0.85 0.74
V-D5 Analysis on Network Output
As explained in Section III-E, the output of the recurrent probabilistic network is a Gaussian mixture distribution to represent the predicted action probabilistic distribution. This is different from previous works that directly regress the actions to execute [18, 46]. This part analyzes the benefits of our probabilistic output.
We first compare the performances of networks based on direct action prediction and probabilistic action prediction. The results are reported in Table VI. We can observe that the recurrent probabilistic network based on probabilistic action prediction outperforms the one using direct action prediction, because the multimodal ambiguity problem is alleviated.
Notably, we argue that the results in Table VI do not sufficiently reveal the value of probabilistic action prediction. This is because the recurrent probabilistic network only observes a single target object due to the object-centric geometry perception block, and the actions in the collected demonstrations are always to manipulate this single object. There is no case like the one shown in Fig. 4, where multiple target objects appear simultaneously in the network observation.
To bridge this gap, we develop a simulated environment as shown in Fig. 18. We use the teleoperation handle shown in Fig. 5 (b) to control the robot hand to collect 50 demonstrations, where we randomly pick up one of these two blocks and put it into the green container. To study multi-target ambiguity, we remove the object-centric geometry perception block in GEM and make the recurrent probabilistic network take images as input. The feature extractor is changed to DINOv2. After training using the collected demonstrations, we compare the success rates of the networks based on direct action prediction and probabilistic action prediction. We find that the one based on direct action prediction makes the robot end-effector hover at the midpoint between the two blocks, leading to a success rate of nearly 0%. By contrast, the network based on probabilistic action prediction achieves a success rate of 87%. These results demonstrate the importance of adopting our proposed probabilistic action prediction.
@@ V-D5 Analysis on Network Output
$\times$ 0.35 0.28 0.20 0.16
$\checkmark$ 0.87 0.78 0.85 0.74
V-D6 Analysis on Hybrid Action Control
The hybrid action control block combines visual servo control and learnable network control to conduct dynamic object manipulation. This experiment studies how this strategy contributes to success rates compared with the common design, i.e., solely based on learnable network control. The performance comparison results between GEMs with and without using the visual servo control are presented in Table VII. We can find that the visual servo control, which converts the original dynamic object manipulation problem into static manipulation, is critical for the overall task completion success rates.
TABLE VIII: Analysis on the effect of simulated demonstrations.
@@ V-D6 Analysis on Hybrid Action Control
$\times$ 0.28 0.45 0.15 0.08
$\checkmark$ 0.92 0.96 0.79 0.74
V-D7 Analysis on Data Efficiency
As explained in Section IV-F, since the recurrent probabilistic network takes uncolored 3D points as input, the observations in simulators become similar to real-world observations. Therefore, the simulated data can directly benefit real-world performances. Correspondingly, we collect 6,600 demonstrations in our built simulation environment. The policy trained using these simulated demonstrations should significantly favor the real-world results. To verify this issue, we compare the real-world performances of the GEM policies with and without being pre-trained with these simulated demonstrations. Both the two policies are tuned using the 80 real-world demonstrations described in Section V-B before the real world evaluation.
The experimental results are presented in Table VIII. We can find that the simulated data boosts the success rates of various skills by large margins. According to the first row of results, only utilizing the 80 real-world demonstrations is insufficient to derive a promising policy, and collecting more real-world demonstrations is labor-intensive. Thankfully, simulated demonstrations can tackle this data scarcity problem because simulated data generation can be efficient.
V-E Generalization Analysis
The various designs in GEM are to improve the generalization of this manipulation system with minimal cost. This subsection conducts experiments to validate the generalization capabilities of GEM from different perspectives.
Figure 19: To study the generalization performance of GEM across different environment backgrounds, we randomly augment the appearance of the conveyor belt, and these generated appearances are unseen in demonstrations. Two augmented samples are visualized in this figure.
@@ V-E Generalization Analysis
Seen 0.88 0.98 0.83 0.80
Unseen 0.91 0.96 0.83 0.81
V-E1 Generalization across Environment Backgrounds
Thanks to the rapid progress of instance segmentation [57, 66], obtaining a segmentation model meeting manipulation precision requirements has become highly cost-effective. Therefore, we design the object-centric geometry perception block in GEM that produces uncolored 3D points appended with instance IDs as the input to the recurrent probabilistic network. In this way, regardless of changes in the environment background, the input to the network remains similar. Therefore, our method should present strong generalization across different environmental backgrounds.
To validate this characteristic, we test the manipulation performance of GEM in different unseen backgrounds by changing the appearance of the conveyor belt. Since changing conveyor belts is expensive to realize in the real world, we conduct this experiment in simulation. As shown in Fig. 19, we randomly augment the colors of the conveyor belt. Since all demonstrations are collected using the conveyor belt illustrated in Fig. 9 (a), these augmented conveyor belts are unseen in demonstrations. The manipulation success rates of GEM in the original seen environment and the augmented unseen environments are presented in Table. IX. We can observe that the performances of GEM are similar in seen and unseen environments, which is consistent with the expectation. In addition, the subsequent sim-to-real experiment in Table XIII indicates that the GEM trained using only simulated data can generalize to real-world manipulation despite the appearance gap between the simulated and real-world environments.
@@ V-E1 Generalization across Environment Backgrounds
Simulated Linear 0.88 0.98 0.83 0.80
Simulated S-curve 0.91 0.97 0.80 0.77
Real Linear 0.92 0.96 0.79 0.74
Real Random curve 0.88 0.98 0.75 0.66
V-E2 Generalization across Motion Patterns
As the visual servo control in the hybrid action control block transforms the original dynamic object manipulation problem into a static manipulation problem, the GEM policy should can generalize to manipulating objects in diverse motion patterns, e.g., various moving velocities and trajectories. The experimental results in Table IV have confirmed that the manipulation success rates remain similar when the conveyor belt speed is adjusted but does not surpass the robot hand tracking velocity limit. In this experiment, we further study this problem by altering the object moving trajectories. Specifically, in simulation, we program the objects to move on the conveyor belt following an S-shaped trajectory as illustrated in Fig. 20 (a). In real-world evaluation, we manually hold a container and put the object to manipulate in this container, moving it along a random curved trajectory, as shown in Fig. 20 (b).
We report the experiment results in Table X. Comparing the performance between linear and curved object movement trajectories, we can conclude that making objects move along an unknown curved trajectory does not harm the manipulation success rates significantly. Therefore, GEM is robust to different motion patterns of moving objects.
@@ V-E2 Generalization across Motion Patterns
TABLE XI: Generalization study on object geometries.
@@
Seen 0.92 0.96 0.79 0.74 0.85
Unseen 0.80 0.90 0.67 0.59 0.74
V-E3 Generalization across Object Geometries
By collecting demonstrations based on objects in diverse geometries and training the recurrent probabilistic network using these demonstrations, GEM can generalize to different geometries. The experiment results in Table I suggest that GEM can perform various manipulation skills on unseen objects with promising success rates. Table II reports the performance of GEM on manipulating seen objects in real-world tasks, where the results are similar to the ones in simulated environments. In this experiment, we further reveal the real-world success rates on manipulating the right nine unseen objects in Fig. 11. The recurrent probabilistic network validated in this experiment is the same as the one in Table II, which is first pre-trained using the 6,600 simulated demonstrations and then fine-tuned with 20 real-world demonstrations per skill.
The experiment results are reported in Table XI. We can find that although the network is only tuned with 20 demonstrations per skill, it still generalizes well to unseen objects in diverse geometries. In Section V-F, we show that if we fine-tune the network using 100 Pick demonstrations, it achieves a success rate of over 97% in an unseen application environment.
V-E4 Generalization across Robot Embodiments
Thanks to the visual servo control, the end-effector moves to a fixed track position relative to the target object before that the recurrent probabilistic network begins generating manipulation actions. Additionally, the robot hand is controlled by the predicted end-effector relative offset based on inverse kinematics. Therefore, GEM is not restricted within a single robot embodiment. To evaluate the robot embodiment generalization, we conduct experiments using four kinds of robot hands in simulated and real-world environments, as shown in Fig. 21. In the simulation, the policy trained using the simulated demonstrations collected from Franka Panda is directly applied to the UR10e robot hand. In the real-world experiment, the policy tuned based on the demonstrations collected with Piper is deployed on the ARX5 robot hand to conduct manipulation.
The experiment results are reported in Table XII. It is observed that employing a robot hand to collect demonstrations and adopting another robot hand to conduct manipulation evaluation achieves similar performance compared with the setting where the demonstration collection and validation are based on the same robot hand. Hence, the robot embodiment generalization capability of GEM is confirmed.

### v2

@@ V-C1 Analysis on Appearance Noise Annealing
Image 0.71 0.35 0.41 0.39
Point 0.93 0.89 0.95 0.87
V-C2 Analysis on Point based Representation
According to Table III, the GEM policy can generalize from simulated data to real-world manipulation to some extent even without the appearance noise annealing strategy. This characteristic is attributed to the 3D point observation of GEM rather than using images. As analyzed in Section III-D, 3D points directly deliver geometry, which is inherently of higher predictivity for estimating manipulation actions. Therefore, by utilizing 3D point observation, GEM naturally prioritizes geometry information over visual appearance, and our developed appearance noise annealing strategy strengthens this prioritization. By contrast, if taking images as input, the network has to infer 3D geometry from visual appearance, meaning that the network must rely on exploiting visual appearance features.
In this experiment, we study the performance gain brought by adopting point observation compared with employing image observation. When the network input is of the image modality, the GEM network encoders are switched to DINOv2 [36], which is a popular image backbone with discriminative representation. The experimental results of adopting point and image observations are presented in Table IV.
According to the results, we can find that the models taking 3D points as input outperform the ones with images significantly. For example, in the zero-shot sim-to-real setting, the point-based policy surpasses the image-based policy by 48%, which confirms the importance of using 3D point observation.
Figure 11: This figure visualizes some instance segmentation cases during manipulation. We render the segmented points in orange and highlight the segmentation mask with an additional orange box.
V-C3 Analysis on Object Segmentation
As depicted in Fig. 6, the GEM network predicts the 3D point segmentation mask of the target object to manipulate, and this segmentation mask is employed to generate tracking actions. Therefore, whether the network can produce a coarse mask of the target object is important for stably tracking the target object.
To enable this stable tracking, we need promising segmentation mask annotations in training data. Thankfully, since we employ simulators to produce demonstrations, we can obtain precise segmentation labels automatically. Besides, since the GEM network primarily utilizes geometry feature to identify the target object, we find the network trained solely using simulation data generalizes to the real world well. To validate this, we collect a real-world segmentation dataset consisting of 1800 point scenes based on the assets shown in Fig. 9. Then, we utilize SAM2 [40], a state-of-the-art open-world segmentation model, to generate segmentation labels efficiently. The obtained labels are employed to validate the GEM network. Following the testing protocol of the COCO benchmark [27], the GEM network achieves a Mask AP of 86.5%, which meets the requirement of stable tracking.
Moreover, we visualize several target object segmentation examples in Fig. 11, where the segmentation mask is highlighted in orange. The results demonstrate promising segmentation performance. As shown in Fig. 11 (a), in the Pick or Rotate task that the operation involves only a single object, the network accurately segments that object. In Fig. 11 (b), for the Put or Insert task that a gripper places an item into a container, the network successfully identifies the item based on its geometry characteristics.
TABLE V: Analysis on the effect of the memory module.
@@ V-C3 Analysis on Object Segmentation
$\times$ 0.86 0.76 0.90 0.82
$\checkmark$ 0.93 0.89 0.95 0.87
V-C4 Analysis on Memory Modeling
In dynamic object manipulation, modeling historical memory enhances the policy’s awareness of both the object’s motion dynamics and task progress. Therefore, we devise the memory module shown in Fig. 6. In experiments, we find that without this module, the robot arm could oscillate at the initial position of manipulation without approaching the target, leading to task failure. This phenomenon is more frequent when a single network learns multiple skills. This is because the policy fails to distinguish whether the manipulation operation has not started or has been completed. Our designed memory module can effectively address this problem by enabling the network to know which actions have been performed. We compare the performances of GEM with and without this memory module and report the results in Table V. According to the results, we can observe that the memory module significantly boosts success rates.
To more thoroughly analyze how long a sequence the memory module can memorize, we design a toy experiment of reciting a random digit sequence. Specifically, we update the structure of this memory module to support accepting a single digit as input and predicting the next digit per step. Then, we randomly generate a sequence of $n$ digits and train the memory module to recite this sequence. After training, we test how many digits can be correctly recited. After repeating this experiment by setting different values of $n$, we find that the memory module can correctly recite a very long sequence like 1000 digits after sufficient epochs of training. Such a memory capability is strong enough for existing embodied policies, because current manipulation tasks usually involve at most tens of times of policy inference.
@@ V-C4 Analysis on Memory Modeling
Direct 0.88 0.80 0.91 0.79
DP 0.81 0.76 0.85 0.63
FM 0.86 0.77 0.88 0.75
Probabilistic 0.93 0.89 0.95 0.87
V-C5 Analysis on Network Output
As explained in Section IV-D, the output of the GEM network is a GMM-based probabilistic distribution to capture the multi-modal nature of actions. This is different from previous works that directly regress the actions to execute [13, 7]. This part analyzes the benefits of adopting this GMM-based probabilistic output.
We first compare the performances of the GEM network outputs based on direct action prediction, diffusion process (DP), flow matching (FM), and GMM. The results are reported in Table VI. We can observe that the GMM-based network outperforms the other counterparts employing the compared output strategies, because the GMM output effectively alleviates the multimodal ambiguity problem. In addition, it is found that the performances of the DP and FM based outputs are inferior to direct action prediction, though DP and FM can also handle the multimodal ambiguity problem. After analysis, we infer this is because the denoising process in DP and FM is challenging for optimization, so a large training batch size is demanded to achieve promising manipulation action generation. In this experiment, since the training demonstration volume is limited, we cannot set such a large batch size. Thus, the obtained success rates are limited. According to these findings, we can summarize that the requirement of a large training batch size makes DP and FM less flexible than our developed GMM-based network output.
Notably, the results in Table VI do not fully demonstrate the advantage of probabilistic action prediction. This is because the objects to manipulate arrive sequentially in dynamic object manipulation on conveyor belts. Hence, at a single moment, the policy usually observes only one object that is suitable for manipulation. As a result, the multi-modal ambiguity that the probabilistic head is designed to resolve remains limited.
To more sufficiently reveal the value of probabilistic action prediction, we develop a simulated environment as shown in Fig. 12. We use the teleoperation handle shown in Fig. 8 (b) to control the robot hand to collect 50 demonstrations, where we randomly pick up one of these two blocks and put it into the green container. After training using the collected demonstrations, we compare the success rates of the networks based on direct action prediction and probabilistic action prediction. We find that the one based on direct action prediction makes the robot end-effector hover at the midpoint between the two blocks, leading to a success rate of nearly 0%. By contrast, the network based on probabilistic action prediction achieves a success rate of 87%. These results demonstrate the importance of adopting our proposed probabilistic action prediction.
@@ V-C5 Analysis on Network Output
$\times$ 0.37 0.31 0.45 0.30
$\checkmark$ 0.93 0.89 0.95 0.87
V-C6 Analysis on Manipulation Action Decomposition
In GEM, we design a strategy that decomposes the manipulation actions into interaction actions and tracking actions, where the interaction actions are decoded from the predicted GMM distribution and the tracking actions are produced based on the point segmentation result. Within this strategy, the tracking actions control the robot end-effector to follow the motion of the target object. In this way, the pose of the end-effector keeps stationary relative to the target object, and thus the interaction actions can be regarded as operating a static object. This design significantly simplifies this task.
The experimental results in Table I and Table II indicate the effectiveness of this strategy. In these two tables, the compared policies marked with an asterisk (*) employ the action decomposition strategy, whereas the ones without an asterisk do not. We can find that the developed strategy improves their success rates by large margins. In addition, we also present the results of GEM without and with the manipulation action decomposition strategy in Table VII. According to the results, it is observed that this strategy is important for the overall performance of GEM.
TABLE VIII: Manipulation performance in various moving speeds.
@@ V-C6 Analysis on Manipulation Action Decomposition
0.05 0.92 1.00 0.91 0.87
0.10 0.94 1.00 0.90 0.86
0.25 0.86 0.98 0.85 0.63
0.50 0.00 0.00 0.00 0.00
V-C7 Analysis on Object Tracking
This work studies dynamic object manipulation on conveyor systems, where the target objects could move at various velocities. As discussed in the above paragraphs, we design a strategy that decomposes the manipulation actions into interaction actions and tracking actions to handle the diverse motion dynamics of target objects. In this part, we validate the performance of GEM on manipulating objects of different velocities.
First of all, we analyze the tracking stability and velocity upper limit in real-world manipulation. Specifically, we adjust the conveyor belt to various speeds and make the robot hand track a moving object based on the visual servo tracking module. It is observed that the maximum stable tracking speed in simulation is approximately 0.25 m/s, and it reaches about 0.3 m/s in the real world. In Fig. 13, we visualize the dynamic curves of the control target signal and corresponding end-effector position. We can find that the end-effector approaches the moving object fast and tracks it stably under the velocity of 0.2 m/s. However, the end-effector fails to track the object when the velocity is 0.6 m/s due to its speed limit.
Subsequently, we validate the success rates of GEM on manipulating objects of different moving speeds in the simulated in-domain experiment setting. The results are reported in Table VIII. Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results, we can observe that the success rates are similar when the conveyor belt runs at a low speed. When the speed is 0.25 m/s, the success rates drop because the robot velocity limit prevents the end-effector from maintaining stable tracking while simultaneously performing precise manipulation. When the speed arrives at 0.50 m/s, the robot hand cannot perform any manipulation task because it cannot keep pace with the movement of the target object.
V-D Generalization Analysis
Thanks to the many designs in GEM, it shows strong generalization in various unseen scenarios, making it suitable for real applications. This subsection analyzes the generalization capabilities of GEM from different perspectives.
Figure 14: To study the background generalization of GEM while excluding the influence of irrelevant factors, we randomly augment the appearance of the conveyor belt in Isaac Gym. The augmented appearances are unseen in demonstrations. Two augmented samples are visualized in this figure.
@@ V-D Generalization Analysis
Seen 0.94 1.00 0.90 0.86
Unseen 0.95 1.00 0.88 0.85
V-D1 Generalization across Environment Backgrounds
In applications, various scenarios show different environment backgrounds. It is expensive to collect demonstrations in every scenario to tune a policy. Thus, we expect that a policy can generalize to unseen scenarios for supporting more flexible deployment. Considering this expectation, we develop GEM, which prioritizes geometry structure and decreases the importance of visual appearance. In this part, we analyze the generalization capability of GEM in unseen backgrounds.
Although we develop the out-of-domain simulation testing protocol, where the policy is trained using demonstrations from Isaac Gym and tested in Isaac Sim, this protocol involves extra influence factors such as simulation parameters. To exclude irrelevant factors, we choose to randomly augment the appearance of the conveyor belt in Isaac Gym, as visualized in Fig. 14. In this way, the policy is trained in Isaac Gym and validated in Isaac Gym, where the environment background is the only difference. We evaluate the performance of GEM and present the results in Table IX.
According to the results, the performances of GEM are similar in environments with seen and unseen backgrounds, which is consistent with the expectation. In addition, the subsequent sim-to-real experiment in Table XIII indicates that the GEM trained using only simulated data can generalize to real-world manipulation despite the appearance gap between the simulated and real-world environments.
@@ V-D1 Generalization across Environment Backgrounds
Simulated Linear 0.94 1.00 0.90 0.86
Simulated S-curve 0.95 1.00 0.88 0.81
Real Linear 0.98 1.00 0.93 0.89
Real Random curve 0.97 1.00 0.90 0.75
V-D2 Generalization across Motion Patterns
As the tracking actions generated based on the manipulation action decomposition strategy transform the original dynamic object manipulation problem into a static object manipulation problem, the GEM policy should be able to generalize to manipulating objects in diverse motion patterns, e.g., various moving velocities and trajectories. The experimental results in Table VIII have confirmed that the manipulation success rates remain similar when the conveyor belt speed is adjusted but does not surpass the robot hand tracking velocity limit. In this experiment, we further study this problem by altering the object moving trajectories. Specifically, in simulation, we program the objects to move on the conveyor belt following an S-shaped trajectory as illustrated in Fig. 15 (a). In real-world evaluation, we manually hold a container and put the object to manipulate in this container, moving it along a random curved trajectory, as shown in Fig. 15 (b).
We report the experiment results in Table X. Comparing the performance between linear and curved object movement trajectories, we can find that making objects move along an unknown curved trajectory does not harm the manipulation success rates significantly, especially for the tasks that do not require precise operation like Put. Therefore, GEM is robust to different motion patterns of moving objects.
@@ V-D2 Generalization across Motion Patterns
TABLE XI: Generalization study on unseen objects.
@@
Seen 0.94 1.00 0.90 0.86 0.93
Unseen 0.89 0.98 0.87 0.85 0.90
V-D3 Generalization across Unseen Objects
Besides environment backgrounds and object motion patterns, generalizing to manipulating unseen objects is also important in real application scenarios. The experimental results reported in Table I and Table II have suggested that the GEM policy can operate novel objects. This advantage is thanks to the training demonstrations collected on objects of diverse geometries, enabling the GEM network to generalize to out-of-domain object geometries. However, there exist extra influence factors in these two experiments, such as inconsistent simulation parameters and new robot embodiments. In this experiment, we further analyze the performance of GEM on manipulating unseen objects after excluding irrelevant factors.
To this end, we import the novel 3D assets shown in Fig. 10 (b) into Isaac Gym to evaluate the GEM trained based on the demonstrations from Isaac Gym. The experiment results are reported in Table XI. We can observe that the success rates of GEM in manipulating out-of-domain objects are similar to the ones of operating in-domain objects. This is because GEM primarily exploits geometry information, and the basic 3D geometries of objects are limited despite their diverse visual appearances. The 3D assets depicted in Fig. 7 (a) have covered many common basic geometries.
V-D4 Generalization across Robot Embodiments
Because of the developed manipulation action decomposition strategy, the end-effector moves to a preset position relative to the target object before the predicted interaction actions are applied. Additionally, the manipulation actions control the joints of a robot based on inverse kinematics. Therefore, GEM is not restricted to a single robot embodiment. To evaluate the robot embodiment generalization, we conduct experiments using four kinds of robot hands in simulated and real-world environments, as shown in Fig. 16. In the simulation based on Isaac Gym, the policy trained using the simulated demonstrations collected from Franka Panda is directly applied to the UR10e robot hand. In the real-robot experiment, the policy tuned based on the demonstrations collected with Piper is deployed on the ARX5 robot hand to conduct manipulation.
The experiment results are reported in Table XII. It is observed that employing a robot to collect demonstrations and adopting another robot to conduct manipulation evaluation achieves similar performance compared with the setting where the demonstration collection and validation are based on the same robot. Hence, the robot embodiment generalization capability of GEM is confirmed.